OpenStudy (anonymous):

The table below shows the fees to check bags on a certain airline. Type of bag Fee per bag First checked bags 20$ Second checked bags 25$ Exra bags bags 50$ If x represents the total number of bags a person checks and each person checks more than two bags, the equation below can be used to determine C, the total fee for all the bags. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all?

OpenStudy (anonymous):

20 + 25 + 50(x - 2) = 145 distribute: 20 + 25 + 50x - 100 = 145 solve for x: -55 + 50x = 145 + 55 50x = 200 /50 x = 4 4 bags were checked.